### v1

@@ 3 The Proofs
  • Accounting for $\mathcal{G}$: In line 3, let $k=\zeta_{t}$ and $(j,j+1)$ be the corresponding max-gap. Suppose that the span $[t,k]$ is capped. The absence of an interval including $j$ and $j+1$ during $[t,k-1]$ implies that $\sum_{t\leq l\leq k}E_{s,l}\leq L+R+m$, where $L$ and $R$ denote the $s$-energy of systems with at most $m-1$ connected components. For reasons we address below, we may assume that $L$ is dominant; hence $R\leq L\leq\mathcal{E}_{m-1,s}$. [^7] The inequality relies on the (easy) fact that the maximum $s$-energy grows monotonically with the number of agents. This is not even needed, however, if we redefine $\mathcal{E}_{m,s}$ as the maximum $s$-energy over all systems with at most $n$ agents and then reason with the value $n^{\prime}\leq n$ that achieves the maximum. It follows that
@@
[…]
    The set $\mathcal{G}$ is a union of spans. If $\zeta_{t}=T_{\delta}$, the last span might not be capped. If so, remove it from $\mathcal{G}$ and call the resulting set $\mathcal{G}^{\prime}$. Summing up, we find that $\sum_{t\in\mathcal{G}^{\prime}}E_{s,t}\leq B\sum_{t\in\mathcal{G}^{\prime}}K_{t}(z)$. If the last span is uncapped then $\zeta_{t}=T_{\delta}$ and no block contains both $j$ and $j+1$ in the span $[t,T_{\delta}]$. The $s$-energy expended in that span is thus of the form $L+R\leq 2\mathcal{E}_{m-1,s}$.
  • Accounting for $\mathcal{U}$: Only ungapped times belong to $\mathcal{U}$, so the 1-energy at time $t\in\mathcal{U}$ is at least $31-\delta-(m_{t}-1)\delta/m\geq 1/$ 3. On the other hand, $E_{s,t}\leq m\leq 3mE_{1,t}\leq 3m\rho K_{t}(z)\leq BK_{t}(z)$.
@@
Actually, the exponent to $1/\rho$ can be reduced to $n-2$, but this is immaterial. By (12),
@@
We begin with the case $m=1$. Assume that $n=2k+1$. At time $t=1$, we have $x_{k}=-x_{-k}=1/2$ and $x_{i}=-x_{-i}=\frac{1}{2}(1-\rho^{i})$, for $0\leq i<k$ and $\rho$ small enough. For $t>1$, we set $x_{i}(t)=(1-\rho^{k})x_{i}(t-1)$. The agents are labeled $-k,\ldots,k$ from left to right. It is easily verified that this constitutes a twist system for the block $[-k,k]$ with initial unit diameter. The $s$-energy $E$ is $\bigl{(}1-\rho^{k}\bigr{)}^{s}E+1$ so, for constant $c>0$,
@@
For the general case, we describe the evolution of an $m$-block twist system with $n$ agents, and denote its $s$-energy by $F(n,m)$: It is assumed that $n-1$ agents are positioned at 00 at time 1 and the last one is at position $1$. If $m=1$, we apply the previous construction after shifting the initial interval from $[-0.5,0.5]$ to $[0,1]$. The initial positions still do not match, but we note that, in a single step, we can move the agents anywhere we want in the interval $[\rho,1-\rho]$ while respecting the constraints of a twist system. This gives us $F(n,1)=1+(1-2\rho)^{s}E$. By adjusting the constant $c$ in (18), the same lower bound still holds.
For $m>1$, at time $1$, we move the agents $n-1$ and $n$ to positions $\rho$ and $1-\rho$, respectively, and we leave the others (if any) at position 0. We then use an $(m-1)$-block twist system recursively for the agents $11,\ldots,n-$ 1. This brings these agents to a common position [^8] To keep the time finite, we can always force completion in a single step once the agents are sufficiently close to each other. in $[0,\rho]$. This gives us the recurrence relation: $F(n,m)\geq 1+\rho^{s}F(n-1,m-1)+(1-2\rho)^{s}F(n,m)$; hence, by induction, for constant $c>0$,
@@
The $s$-energy is often used for small $s$, so we state the case of $s=O\bigl{(}1/m\log\frac{1}{\rho}\bigr{)}$, which matches the bound of Theorem 1.1 (a) for $m=n-1$.
@@
.absent\!\!\!.\,\,. $\mathcal{E}_{m,s}^{\text{\sc tw}}\geq(c/s)^{m}(1/\rho)^{\lfloor(n+m-1)/2\rfloor}$, for constant $c>0$, small enough $\rho$ and $s=O\bigl{(}1/m\log\frac{1}{\rho}\bigr{)}$.
@@
.absent\!\!\!.\,\,. If $G_{\leq t_{c}}$ is connected, then $\sum_{t\leq t_{c}}D_{t}\geq\,\rho n^{-2}\,\|x-\hat{x}\,\|_{q}^{2}\,$.
Proof. Let $G_{t_{0}}$ denote the graph over $n$ vertices with no edges. We define $t_{1},\ldots,t_{c}$ as the sequence of times $t$ at which the addition of $G_{t}$ reduces the number of connected components in $G_{\leq t-1}$. At any time $t_{k}$ ($k>0$), the drop $d_{k}$ in the number of components can be achieved by $d_{k}$ edges from $G_{t_{k}}$. Let $F_{k}$ denote such a set of edges: we can always order $F_{k}$ so that every edge in the sequence contains at least one vertex not encountered yet. This shows that the sum of the squared lengths of the edges in $F_{k}$ does not exceed $D_{t_{k}}$. We note that $F:=F_{1}\cup\cdots\cup F_{c}$ forms a collection of $n-1$ edges from (the connected graph) $G_{\leq t_{c}}$ and $F$ spans all $n$ vertices.
Consider the intervals formed by the edges in $F_{k}$ at time $t_{k}$, for all $k\in[c]$. The union of these intervals covers the smallest interval $[a,b]$ enclosing all the vertices at time 00 (and hence at all times). To see why, pick any $z$ such that $a<z<b$ and denote by $L$ and $R$ the vertices on both sides of $z$ at time 00. Neither set is empty and, by convexity, both of them remain on their respective side of $z$ until an edge of some $G_{t}$ joins $L$ to $R$. When that happens (which it must since $G_{\leq t_{c}}$ is connected), the joining edge(s) reduce(s) the number of components of $G_{\leq t-1}$ by at least one, so $F$ must grab at least one of them, which proves our claim. Let $l_{1},\ldots,l_{n-1}$ denote the lengths of the edges of $F$ (at the time of their insertion). By Cauchy-Schwarz, $\sum_{t\leq t_{c}}D_{t}\geq\sum_{i=1}^{n-1}l_{i}^{2}\geq(b-a)^{2}/(n-1)$. The lemma follows from the inequalities $\|x-\hat{x}\,\|_{q}^{2}\leq\|q\|_{1}(b-a)^{2}\leq(n/\rho)(b-a)^{2}$. $\Box$
@@
For the general case, we denote by $F(n,m)$ the $s$-energy of the system with initial diameter equal to 1. We showed that $F(n,1)\geq bn^{2-s}/\rho s$. We now describe the steps of the dynamics for $m>1$. To simplify the notation, we assume that $\nu:=n/m$ is an integer. [^10] This can be relaxed with a simple padding argument we may omit. For $i\in[m]$, let $C_{i}$ be the path linking vertices $[(i-1)\nu+1,i\nu]$.
@@
    At time $t=1$, the vertices of $C_{1}$ are placed at position 00 while all the others are stationed at $1$. The paths $C_{1}$ and $C_{2}$ are linked together into a single path so the system has $m-1$ components. Vertices $\nu$ and $\nu+1$ move to positions $\rho$ and $1-\rho$ respectively while the others do not move at all. The $s$-energy expended during that step is equal to 1.
@@
.absent\!\!\!.\,\,. There exist reversible averaging systems with initial diameter equal to 1 whose $s$-energy is at least $\bigl{(}c/s\rho^{1-s}\bigr{)}^{m}(n/m)^{(1-s)m+1}$, for constant $c>0$. The number of vertices is $n$ and the number of connected components is bounded by $m$; furthermore, all positive entries in the stochastic matrices are at least $\rho$, with $40<\rho<1/$ 4.
@@
.absent\!\!\!.\,\,. If $G=(V,E)$ is connected, then $\sum_{(i,j)\in E}\,(x_{i}-x_{j})^{2}\geq b(h/d)^{2}\,\|x-\hat{x}\|_{q}^{2}\geq b(h\Delta)^{2}/2d$, for constant $b>0$, where $\Delta$ is the diameter of the agent positions $x_{1},\ldots,x_{n}$.
Proof. All of the ideas in this proof come from [4, 58]. The inequality is invariant under shifting and scaling, so we may assume that $\hat{x}=\mathbf{0}$ and $\|x\|_{2}=1$. Relabel the coordinates of $x$ so they appear in nonincreasing order, and define $y\in\mathbb{R}^{n}$ such that $y_{i}=\max\{x_{i},0\}$. Let $\alpha=\mathrm{argmax}_{k}(y_{k}>0)$ and $\beta=\min\{\alpha,\lfloor n/2\rfloor\}$. By switching $x$ into $-x$ if necessary, [^12] Intuitively, by changing all signs if necessary, we force the minority sign among the coordinates of $x$ to be positive unless their contribution to the norm of $x$ is too small. we can always assume that $\|y\|_{2}^{2}>c\!:=1/6$ if $\alpha=\beta$, and $\|y\|_{2}^{2}\geq 1-c$ if $\alpha>\beta$. By Cauchy-Schwarz, $(y_{i}+y_{j})^{2}\leq 2(y_{i}^{2}+y_{j}^{2})$; hence,
@@
Suppose that $\alpha>\beta=\lfloor n/2\rfloor$. It follows from $\sum_{i=1}^{n}x_{i}=0$ that $\sum_{i=\alpha+1}^{n}|x_{i}|=\|y\|_{1}\geq(\beta+1)y_{\beta+1}$. By Cauchy-Schwarz, this yields $ny_{\beta+1}^{2}/4\leq\sum_{i=\alpha+1}^{n}x_{i}^{2}=1-\|y\|_{2}^{2}$, and by (24), $\sum_{(i,j)\in E}\,\bigl{|}\,y_{i}^{2}-y_{j}^{2}\,\bigr{|}\geq(1-5c)h=ch$. If $\alpha=\beta$, then $y_{\lfloor n/2\rfloor+1}=0$ and $\sum_{(i,j)\in E}\,\bigl{|}\,y_{i}^{2}-y_{j}^{2}\,\bigr{|}\geq h\|y\|_{2}^{2}>ch$. Applying (23) shows that $\sum_{(i,j)\in E}(x_{i}-x_{j})^{2}\geq(ch)^{2}/2d$, which gives us the first inequality of the lemma. The second one follows from the fact that the interval $[u,v]$ enclosing the vertex positions contains 00. By Cauchy-Schwarz, $21=\|x\|_{2}^{2}\geq u^{2}+v^{2}\geq\frac{1}{2}(v+|u|)^{2}=\Delta^{2}/$ 2, and the proof is complete. $\Box$
@@
.absent\!\!\!.\,\,. If $G_{\leq t_{c}}$ is connected, then $\sum_{t\leq t_{c}}\sum_{i=1}^{m_{t}}\Delta_{t,i}^{2}\geq\,\frac{1}{2dmn}\|x-\hat{x}\,\|_{q}^{2}\,$.
Proof. At any time $t_{k}>0$, the drop $d_{k}$ in the number of components can be achieved by $d_{k}$ (or fewer) components in $G_{t_{k}}$. We collect the intervals spanned by these components into a set $F$, to which we add the intervals for the components of $G_{0}$; thus $|F|<2m$. A simple convexity argument (omitted) shows that the union of the intervals in $F$ coincides with the interval $[a,b]$ enclosing the $n$ vertices at time $1$; so the lengths $l_{1},\ldots,l_{|F|}$ of the intervals in $F$ sum up to at least $b-a$. By Cauchy-Schwarz, $\sum_{t\leq t_{c}}\sum_{i=1}^{m_{t}}\Delta_{t,i}^{2}\geq\sum_{i=1}^{|F|}l_{i}^{2}\geq(b-a)^{2}/(2m-1)$. The lemma follows from $\|x-\hat{x}\,\|_{q}^{2}\leq\|q\|_{1}(b-a)^{2}\leq dn(b-a)^{2}$. $\Box$
@@
It is assumed here that each graph $G_{t}$ is picked independently, uniformly from the set of simple $(d-1)$-regular graphs with $n$ vertices, with fixed $d>3$ [66]. (We use $d-1$ because $d$ must account for the self-loops.) The system is a special case of a rev, so we use the same notation. The stochastic matrix $P_{t}$ for $G_{t}$ is $\frac{1}{d}M_{t}$, where $M_{t}$ is a random symmetric 0/1 matrix with a positive diagonal and all row sums equal to $d$; we have $q=d\mathbf{1}$. Recall that the $s$-energy $\mathcal{E}_{s}^{\text{\sc ran}}$ is defined over systems with unit variance $\|x-\hat{x}\,\|_{q}^{2}=1$, where $\hat{x}=\frac{1}{d}\langle x,\mathbf{1}\rangle$. Let $x_{i},y_{i}$ be the positions of agent $i$ at step $t$ and $t+1$ respectively. As usual, we may place the center of gravity $\hat{x}$ at the origin at time 0, where it will remain forever; that is, $\sum_{i=1}^{n}x_{i}=0$.
@@
.absent\!\!\!.\,\,. $\mathbb{E}\sum_{i=1}^{n}y_{i}^{2}=(1-b)\sum_{i=1}^{n}x_{i}^{2}$, where $b=\frac{(d-1)n}{d(n-1)}\,$.
@@
$\begin{split}\sum_{i}x_{i}^{2}-\sum_{i}y_{i}^{2}&=\sum_{i}x_{i}^{2}-\sum_{i}\left(x_{i}-\frac{1}{d}\sum_{j}m_{ij}\delta_{ij}\right)^{2}\\ &=\frac{2}{d}\sum_{i,j}m_{ij}x_{i}\delta_{ij}-\frac{1}{d^{2}}\sum_{i,j,k}m_{ij}m_{ik}\delta_{ij}\delta_{ik}\\ &=\frac{1}{d}\sum_{i,j}m_{ij}\delta_{ij}^{2}-\frac{1}{2d^{2}}\sum_{i,j,k}m_{ij}m_{ik}(\delta_{ij}^{2}+\delta_{ik}^{2}-\delta_{jk}^{2})\\ &=\frac{1}{2d^{2}}\sum_{i,j,k:i\neq j}m_{ik}m_{jk}\delta_{ij}^{2}\,,\end{split}$ (27)

### v2

@@ 3 Energy Bounds for Averaging Systems
  • Accounting for $\mathcal{G}$: In line 3, let $k=\zeta_{t}$ and $(j,j+1)$ be the corresponding max-gap. Suppose that the span $[t,k]$ is capped. The absence of an interval including $j$ and $j+1$ during [t,k1][t,k-1] implies that $\sum_{t\leq l\leq k}E_{s,l}\leq L+R+m$, where $L$ and $R$ denote the $s$-energy of systems with at most m1m-1 connected components. For reasons we address below, we may assume that $L$ is dominant; hence $R\leq L\leq\mathcal{E}_{m-1,s}$. [^5] The inequality relies on the (easy) fact that the maximum $s$-energy grows monotonically with the number of agents. This is not even needed, however, if we redefine $\mathcal{E}_{m,s}$ as the maximum $s$-energy over all systems with at most $n$ agents and then reason with the value $n^{\prime}\leq n$ that achieves the maximum. It follows that
@@
    $\sum_{l=t}^{k}E_{s,l}\leq B\bigl(x_{v}(k)-x_{j}(k)\bigr)z^{j},\hskip 11.38092pt\text{\small{with}}\hskip 5.69046ptB=9c_{m-1}m\rho(1/s)^{m-1}$ (7)
    The set $\mathcal{G}$ is a union of spans. If $\zeta_{t}=T_{\delta}$, the last span might not be capped. If so, remove it from $\mathcal{G}$ and call the resulting set $\mathcal{G}^{\prime}$. Summing up, we find that $\sum_{t\in\mathcal{G}^{\prime}}E_{s,t}\leq B\sum_{t\in\mathcal{G}^{\prime}}K_{t}(z)$. If the last span is uncapped then $\zeta_{t}=T_{\delta}$ and no block contains both $j$ and $j+1$ in the span $t,T_{\delta}]$. The $s$-energy expended in that span is thus of the form $L+R\leq 2\mathcal{E}_{m-1,s}$.
  • Accounting for $\mathcal{U}$: Only ungapped times belong to $\mathcal{U}$, so the 1-energy at time $t\in\mathcal{U}$ is at least $1-\delta-(m_{t}-1)\delta/m\geq 1/3$. On the other hand, $E_{s,t}\leq m\leq 3mE_{1,t}\leq 3m\rho K_{t}(z)\leq BK_{t}(z)$.
@@
Actually, the exponent to $1/\rho$ can be reduced to n2n-2, but this is immaterial. By (2),
@@
We begin with the case $m=1$. Assume that $n=2k+1$. At time $t=1$, we have $x_{k}=-x_{-k}=1/2$ and $x_{i}=-x_{-i}=\frac{1}{2}(1-\rho^{i})$, for $0\leq i<k$ and $\rho$ small enough. For $t>1$, we set $x_{i}(t)=(1-\rho^{k})x_{i}(t-1)$. The agents are labeled $k,\ldots,k$ from left to right. It is easily verified that this constitutes a twist system for the block [k,k][-k,k] with initial unit diameter. The $s$-energy $E$ is $\bigl(1-\rho^{k}\bigr)^{s}E+1$ so, for constant $c>0$,
@@
For the general case, we describe the evolution of an $m$-block twist system with $n$ agents, and denote its $s$-energy by $F(n,m)$: It is assumed that n1n-1 agents are positioned at 0 at time 1 and the last one is at position $1$. If $m=1$, we apply the previous construction after shifting the initial interval from [0.5,0.5][-0.5,0.5] to $[0,1]$. The initial positions still do not match, but we note that, in a single step, we can move the agents anywhere we want in the interval $\rho,1-\rho]$ while respecting the constraints of a twist system. This gives us $F(n,1)=1+(1-2\rho)^{s}E$. By adjusting the constant $c$ in (8), the same lower bound still holds.
For $m>1$, at time $1$, we move the agents n1n-1 and $n$ to positions $\rho$ and $1-\rho$, respectively, and we leave the others (if any) at position 0. We then use an (m1)(m-1)-block twist system recursively for the agents $1,\ldots,n-1$. This brings these agents to a common position [^6] To keep the time finite, we can always force completion in a single step once the agents are sufficiently close to each other. in $0,\rho]$. This gives us the recurrence relation: $F(n,m)\geq 1+\rho^{s}F(n-1,m-1)+(1-2\rho)^{s}F(n,m)$; hence, by induction, for constant $c>0$,
@@
The $s$-energy is often used for small $s$, so we state the case of $s=O\bigl(1/m\log\frac{1}{\rho}\bigr)$, which matches the bound of Theorem 2.1 (a) for m=n1m=n-1.
@@
.\!\!\!.\,\, $\mathcal{E}_{m,s}^{\text{\sc tw}}\geq(c/s)^{m}(1/\rho)^{\lfloor(n+m-1)/2\rfloor}$, for constant $c>0$, small enough $\rho$ and $s=O\bigl(1/m\log\frac{1}{\rho}\bigr)$.
@@
.\!\!\!.\,\, If $G_{\leq t_{c}}$ is connected, then $\sum_{t\leq t_{c}}D_{t}\geq\,\rho n^{-2}\,\|x-\hat{x}\,\|_{q}^{2}\,$.
Proof. Let $G_{t_{0}}$ denote the graph over $n$ vertices with no edges. We define $t_{1},\ldots,t_{c}$ as the sequence of times $t$ at which the addition of $G_{t}$ reduces the number of connected components in $G_{\leq t-1}$. At any time $t_{k}$ ($k>0$), the drop $d_{k}$ in the number of components can be achieved by $d_{k}$ edges from $G_{t_{k}}$. Let $F_{k}$ denote such a set of edges: we can always order $F_{k}$ so that every edge in the sequence contains at least one vertex not encountered yet. This shows that the sum of the squared lengths of the edges in $F_{k}$ does not exceed $D_{t_{k}}$. We note that $F:=F_{1}\cup\cdots\cup F_{c}$ forms a collection of n1n-1 edges from (the connected graph) $G_{\leq t_{c}}$ and $F$ spans all $n$ vertices.
Consider the intervals formed by the edges in $F_{k}$ at time $t_{k}$, for all $k\in[c]$. The union of these intervals covers the smallest interval $[a,b]$ enclosing all the vertices at time 0 (and hence at all times). To see why, pick any $z$ such that $a<z<b$ and denote by $L$ and $R$ the vertices on both sides of $z$ at time 0. Neither set is empty and, by convexity, both of them remain on their respective side of $z$ until an edge of some $G_{t}$ joins $L$ to $R$. When that happens (which it must since $G_{\leq t_{c}}$ is connected), the joining edge(s) reduce(s) the number of components of $G_{\leq t-1}$ by at least one, so $F$ must grab at least one of them, which proves our claim. Let $l_{1},\ldots,l_{n-1}$ denote the lengths of the edges of $F$ (at the time of their insertion). By Cauchy-Schwarz, $\sum_{t\leq t_{c}}D_{t}\geq\sum_{i=1}^{n-1}l_{i}^{2}\geq(b-a)^{2}/(n-1)$. The lemma follows from the inequalities $\|x-\hat{x}\,\|_{q}^{2}\leq\|q\|_{1}(b-a)^{2}\leq(n/\rho)(b-a)^{2}$. $\Box$
@@
For the general case, we denote by $F(n,m)$ the $s$-energy of the system with initial diameter equal to 1. We showed that $F(n,1)\geq bn^{2-s}/\rho s$. We now describe the steps of the dynamics for $m>1$. To simplify the notation, we assume that $\nu:=n/m$ is an integer. [^8] This can be relaxed with a simple padding argument we may omit. For $i\in[m]$, let $C_{i}$ be the path linking vertices $(i-1)\nu+1,i\nu]$.
@@
    At time $t=1$, the vertices of $C_{1}$ are placed at position 0 while all the others are stationed at $1$. The paths $C_{1}$ and $C_{2}$ are linked together into a single path so the system has m1m-1 components. Vertices $\nu$ and $\nu+1$ move to positions $\rho$ and $1-\rho$ respectively while the others do not move at all. The $s$-energy expended during that step is equal to 1.
@@
.\!\!\!.\,\, There exist reversible averaging systems with initial diameter equal to 1 whose $s$-energy is at least $\bigl(c/s\rho^{1-s}\bigr)^{m}(n/m)^{(1-s)m+1}$, for constant $c>0$. The number of vertices is $n$ and the number of connected components is bounded by $m$; furthermore, all positive entries in the stochastic matrices are at least $\rho$, with $0<\rho<1/4$.
@@
.\!\!\!.\,\, If $G=(V,E)$ is connected, then $\sum_{(i,j)\in E}\,(x_{i}-x_{j})^{2}\geq b(h/d)^{2}\,\|x-\hat{x}\|_{q}^{2}\geq b(h\Delta)^{2}/2d$, for constant $b>0$, where $\Delta$ is the diameter of the agent positions $x_{1},\ldots,x_{n}$.
Proof. All of the ideas in this proof come from [4, 58]. The inequality is invariant under shifting and scaling, so we may assume that $\hat{x}=\mathbf{0}$ and $\|x\|_{2}=1$. Relabel the coordinates of $x$ so they appear in nonincreasing order, and define $y\in\mathbb{R}^{n}$ such that $y_{i}=\max\{x_{i},0\}$. Let $\alpha=\mathrm{argmax}_{k}(y_{k}>0)$ and $\beta=\min\{\alpha,\lfloor n/2\rfloor\}$. By switching $x$ into x-x if necessary, [^10] Intuitively, by changing all signs if necessary, we force the minority sign among the coordinates of $x$ to be positive unless their contribution to the norm of $x$ is too small. we can always assume that $\|y\|_{2}^{2}>c\!:=1/6$ if $\alpha=\beta$, and $\|y\|_{2}^{2}\geq 1-c$ if $\alpha>\beta$. By Cauchy-Schwarz, $(y_{i}+y_{j})^{2}\leq 2(y_{i}^{2}+y_{j}^{2})$; hence,
@@
Suppose that $\alpha>\beta=\lfloor n/2\rfloor$. It follows from $\sum_{i=1}^{n}x_{i}=0$ that $\sum_{i=\alpha+1}^{n}|x_{i}|=\|y\|_{1}\geq(\beta+1)y_{\beta+1}$. By Cauchy-Schwarz, this yields $ny_{\beta+1}^{2}/4\leq\sum_{i=\alpha+1}^{n}x_{i}^{2}=1-\|y\|_{2}^{2}$, and by (14), $\sum_{(i,j)\in E}\,\bigl|\,y_{i}^{2}-y_{j}^{2}\,\bigr|\geq(1-5c)h=ch$. If $\alpha=\beta$, then $y_{\lfloor n/2\rfloor+1}=0$ and $\sum_{(i,j)\in E}\,\bigl|\,y_{i}^{2}-y_{j}^{2}\,\bigr|\geq h\|y\|_{2}^{2}>ch$. Applying (13) shows that $\sum_{(i,j)\in E}(x_{i}-x_{j})^{2}\geq(ch)^{2}/2d$, which gives us the first inequality of the lemma. The second one follows from the fact that the interval $[u,v]$ enclosing the vertex positions contains 0. By Cauchy-Schwarz, $1=\|x\|_{2}^{2}\geq u^{2}+v^{2}\geq\frac{1}{2}(v+|u|)^{2}=\Delta^{2}/2$, and the proof is complete. $\Box$
@@
.\!\!\!.\,\, If $G_{\leq t_{c}}$ is connected, then $\sum_{t\leq t_{c}}\sum_{i=1}^{m_{t}}\Delta_{t,i}^{2}\geq\,\frac{1}{2dmn}\|x-\hat{x}\,\|_{q}^{2}\,$.
Proof. At any time $t_{k}>0$, the drop $d_{k}$ in the number of components can be achieved by $d_{k}$ (or fewer) components in $G_{t_{k}}$. We collect the intervals spanned by these components into a set $F$, to which we add the intervals for the components of $G_{0}$; thus $|F|<2m$. A simple convexity argument (omitted) shows that the union of the intervals in $F$ coincides with the interval $[a,b]$ enclosing the $n$ vertices at time $1$; so the lengths $l_{1},\ldots,l_{|F|}$ of the intervals in $F$ sum up to at least bab-a. By Cauchy-Schwarz, $\sum_{t\leq t_{c}}\sum_{i=1}^{m_{t}}\Delta_{t,i}^{2}\geq\sum_{i=1}^{|F|}l_{i}^{2}\geq(b-a)^{2}/(2m-1)$. The lemma follows from $\|x-\hat{x}\,\|_{q}^{2}\leq\|q\|_{1}(b-a)^{2}\leq dn(b-a)^{2}$. $\Box$
@@
It is assumed here that each graph $G_{t}$ is picked independently, uniformly from the set of simple (d1)(d-1)-regular graphs with $n$ vertices, with fixed $d>3$ [66]. (We use d1d-1 because $d$ must account for the self-loops.) The system is a special case of a rev, so we use the same notation. The stochastic matrix $P_{t}$ for $G_{t}$ is $\frac{1}{d}M_{t}$, where $M_{t}$ is a random symmetric 0/1 matrix with a positive diagonal and all row sums equal to $d$; we have $q=d\mathbf{1}$. Recall that the $s$-energy $\mathcal{E}_{s}^{\text{\sc ran}}$ is defined over systems with unit variance $\|x-\hat{x}\,\|_{q}^{2}=1$, where $\hat{x}=\frac{1}{d}\langle x,\mathbf{1}\rangle$. Let $x_{i},y_{i}$ be the positions of agent $i$ at step $t$ and $t+1$ respectively. As usual, we may place the center of gravity $\hat{x}$ at the origin at time 0, where it will remain forever; that is, $\sum_{i=1}^{n}x_{i}=0$.
@@
.\!\!\!.\,\, $\mathbb{E}\sum_{i=1}^{n}y_{i}^{2}=(1-b)\sum_{i=1}^{n}x_{i}^{2}$, where $b=\frac{(d-1)n}{d(n-1)}\,$.
@@
[…]